OpenStudy (anonymous):

find the slope, if it exists, ofcontaining the pair of points (3,1) and (5,-3)

OpenStudy (anonymous):

rise/run y2-y1 / x2 - x1 m=-2 -3 -1 = 4 5 - 3 = 2 4/2 = 2

OpenStudy (anonymous):

Thanks but I'm lost the final answer is

OpenStudy (anonymous):

there is a negative in there ...i'm famous for dropping them (x,y) is a point ya? in (3,1) the 3 is an x and the 1 is the y.... that is the first ones so call them x1 and y1 in (5,-3) the 5 is the x and -3 is the y .... that is the second set so call them x2 and y2 rise is y2 - y1 -3 -1 = -4 run is x2 - x1 5 - 3 = 2 -4/2 = -2/1 = -2

OpenStudy (anonymous):

graphing it makes it even easier: draw the two points and then draw a line connecting them. If the line points down to your right it is a negative slope. If it points up to the right then it is a positive slope

OpenStudy (anonymous):

thank you